OpenStudy (anonymous):

solve by using power series method: 2y'-y=cosh(x). compute the first 6 coefficients (a0-a5)

OpenStudy (primeralph):

Okay, that's a little long. What part of it don't you get?

OpenStudy (anonymous):

yeah, i have doubt in one part... I'll show you

OpenStudy (anonymous):

Assume that y has power series y= ∑_(n=0)^∞▒〖a_n x^n 〗 Then y^'= ∑_(n=0)^∞▒〖na_n x^(n-1) 〗 so that 2y^'-y=∑_(n=0)^∞▒〖2na_n x^(n-1)-a_n x^n 〗 Collect the above series with powers of x. ∑_(n=0)^(∞ )▒〖 █(2na_n x^(n-1)-a_n x^n=2(a_1+2a_2 x+3a_3 x^2+4a_4 x^3+5a_5 x^4…)-@(a_0+a_1 x+a_2 x^2+a_3 x^3+a_4 x^4+a_5 x^5…)@)〗 =(2a_1-a_0 )+(2∙2a_2-a_1 )x+(2∙3a_3-a_2 ) x^2+(2∙4a_4-a_3 ) x^3+⋯ After that, I'm confused I don't know if is: cosh⁡x=1+x^2/2!+x^4/4!+x^6/6!+⋯ 〗 or cosh⁡x=1+x^2/2+x^4/6+x^6/24+x^8/120⋯ 〗 I think im having problems with the algebra

OpenStudy (primeralph):

I'll just give you the final answer.

OpenStudy (primeralph):

..........because I can't see the stuff you typed up......

OpenStudy (anonymous):

Have you seen my other post on the almost identical problem?

OpenStudy (anonymous):

with the senh(x)?

OpenStudy (anonymous):

Do you know the power series for \(\cos x\)?$$\cos x=\sum_{n=0}^\infty\frac{(-1)^n}{(2n)!}x^{2n}$$

OpenStudy (primeralph):